For all questions: (a) state the null and alternative hypotheses, (b) conduct appropriate statistical methods with alpha-.05, (c) report your final statistics with a decision regarding the null hypothesis, and (d) summarize your results in your own words. 1. In this era of credit cards there is discussion on whether or not coin use is declining. A professor has been collecting data for years on the number of cents people carry in their pockets or purses and maintains that the population mean is 2.2 cents. She polls her current class and obtains the following data: 1, 2, 3, 0, 0, 19, 43, 2, 7, 9, 0, 0, 1, 0, 5. Has there been a change in the number of cents people carry? (Report the appropriate statistics and state whether you reject or accept the null hypothesis)
2. Informal interviews indicated that most people would leave a parking lot faster if a car were waiting for them to depart. To verify this, 30 drivers were observed and timed for how long it took each one to enter his or her car and leave the parking space. Fifteen were randomly selected to be intruded upon by having another car pull up near the parking space as if its driver were waiting to claim the vacated space. The other 15 were timed but not intruded upon. The data (in seconds) are given here. Does driver behavior confirm what the informal interviews indicated? (Report the t statistic and state whether you reject or accept the null hypothesis) Intruded-Upon Drivers: 51. 58. 46. 80. 83. 45. 63. 69. 63. 69. 65. 58. 75. 64. 63 SD-11.06 Variance-122.41 Nonintruded-Upon Drivers: 61. 55. 46. 70. 52. 55. 68. 53. 52. 62. 45. 55. 65. 41. 52 SD-8.38 Variance-70.27
3. In a study of the effect of violent material on aggression, researchers test the aggressiveness of 9 volunteers. First, t searchers test the aggressiveness of t expose them to a violent film. Afterwards, the researchers retest the volunteers aggressiveness. Based on the data listed below what do you conclude? he volunteers. Next they 4 10 4 4 4. Twenty five high school students complete a preparation program for taking the SAT test. Here are the SAT scores fronm the 25 students who completed the SATp rep program 434 694 457 534 720 400 484 478 610 641 425 636 454 514 563 370 499 640 501 625 612 471 598 509 531 The mean of these scores is 536.00. We know that the population average for SAT scores is 500 with a standard deviation of 100 The question is, are these students SAT scores (M 536) significantly greater than a population mean of 500 with a population standard deviation of 100? Note that the creators of the SAT prep program claims that it will increase (and not decrease) your SAT score
Answer #1

(1)

Data:    

n = 15   

μ = 2.2   

s = 11.40092   

x-bar = 6.13333   

Hypotheses:   

Ho: μ = 2.2   

Ha: μ ≠ 2.2   

Decision Rule:   

α = 0.05   

Degrees of freedom = 15 - 1 = 14

Lower Critical t- score = -2.144786681

Upper Critical t- score = 2.144786681

Reject Ho if |t| > 2.144786681

Test Statistic:   

SE = s/√n = 11.40092/√15 = 2.943704886

t = (x-bar - μ)/SE = (6.13333 - 2.2)/2.9437048860962 = 1.336184

p- value = 0.202802752   

Decision (in terms of the hypotheses):

Since 1.336183535 < 2.144787 we fail to reject Ho

Conclusion (in terms of the problem):

There is no sufficient evidence that there is a change in the number of cents people carry.